# EUROPEAN ORGANIZATION FOR NUCLEAR RESEARCH (CERN) CERN-EP-2023-280 LHCb-PAPER-2023-037

### v1

@@ I. Kostiuk
A. Kozachuk
P. Kravchenko
L. Kravchuk
M. Kreps
S. Kretzschmar
P. Krokovny
W. Krupa
W. Krzemien
J. Kubat,
S. Kubis
W. Kucewicz
M. Kucharczyk
V. Kudryavtsev
E. Kulikova
A. Kupsc
B. K. Kutsenko
D. Lacarrere
A. Lai
A. Lampis
D. Lancierini
C. Landesa Gomez
J.J. Lane
R. Lane
C. Langenbruch
J. Langer
O. Lantwin
T. Latham
F. Lazzari
C. Lazzeroni
R. Le Gac
S.H. Lee
R. Lefèvre
A. Leflat
S. Legotin
M. Lehuraux
O. Leroy
T. Lesiak
B. Leverington
A. Li
H. Li
K. Li
L. Li
P. Li
P.-R. Li
S. Li
T. Li
T. Li
Y. Li,
Y. Li
Z. Li
Z. Lian
X. Liang
C. Lin
T. Lin
R. Lindner
V. Lisovskyi
R. Litvinov
F. L. Liu
G. Liu
K. Liu
Q. Liu
S. Liu
Y. Liu
Y. Liu,
Y. L. Liu
A. Lobo Salvia
A. Loi
J. Lomba Castro
T. Long
J.H. Lopes
A. Lopez Huertas
S. López Soliño
G.H. Lovell
C. Lucarelli
D. Lucchesi
S. Luchuk
M. Lucio Martinez
V. Lukashenko
Y. Luo
A. Lupato
E. Luppi
K. Lynch
X.-R. Lyu
G. M. Ma
R. Ma
S. Maccolini
F. Machefert
F. Maciuc
B. M. Mack
I. Mackay
L. M. Mackey
L.R. Madhan Mohan
M. M. Madurai
A. Maevskiy
D. Magdalinski
V. Mahajan
D. Maisuzenko
M.W. Majewski,
J.J. Malczewski
S. Malde
B. Malecki
L. Malentacca,
A. Malinin
T. Maltsev
G. Manca
G. Mancinelli
C. Mancuso
R. Manera Escalero,
D. Manuzzi
D. Marangotto
J.F. Marchand
R. Marchevski
U. Marconi
S. Mariani
C. Marin Benito
J. Marks
A.M. Marshall
P.J. Marshall,
G. Martelli
G. Martellotti
L. Martinazzoli
M. Martinelli
D. Martinez Santos
F. Martinez Vidal
A. Massafferri
M. Materok
R. Matev
A. Mathad
V. Matiunin
C. Matteuzzi
K.R. Mattioli
A. Mauri
E. Maurice
J. Mauricio
P. Mayencourt
M. Mazurek
M. McCann
L. Mcconnell
T.H. McGrath
N.T. McHugh
A. McNab
R. McNulty
B. Meadows
G. Meier
D. Melnychuk
M. Merk
A. Merli
L. Meyer Garcia
D. Miao
H. Miao
M. Mikhasenko
D.A. Milanes
A. Minotti
E. Minucci
T. Miralles
S.E. Mitchell
B. Mitreska
D.S. Mitzel
A. Modak
A. Mödden
R.A. Mohammed
R.D. Moise
S. Mokhnenko
T. Mombächer
M. Monk
I.A. Monroy
S. Monteil
A. Morcillo Gomez
G. Morello
M.J. Morello
M.P. Morgenthaler
A.B. Morris
A.G. Morris
R. Mountain
H. Mu
Z. M. Mu
E. Muhammad
F. Muheim
M. Mulder
K. Müller
F. Muñoz-Rojas
R. Murta
P. Naik
T. Nakada
R. Nandakumar
T. Nanut
I. Nasteva
M. Needham
N. Neri
S. Neubert
N. Neufeld
P. Neustroev,
R. Newcombe,
J. Nicolini
D. Nicotra
E.M. Niel
N. Nikitin
P. Nogga,
N.S. Nolte
C. Normand
J. Novoa Fernandez
G. Nowak
C. Nunez
H. N. Nur
A. Oblakowska-Mucha
V. Obraztsov
T. Oeser
S. Okamura
R. Oldeman
F. Oliva
M. Olocco
C.J.G. Onderwater
R.H. O’Neil
J.M. Otalora Goicochea
T. Ovsiannikova
P. Owen
A. Oyanguren
O. Ozcelik
K.O. Padeken
B. Pagare
P.R. Pais
T. Pajero
A. Palano
M. Palutan
G. Panshin
L. Paolucci
A. Papanestis
M. Pappagallo
L.L. Pappalardo
C. Pappenheimer
C. Parkes
B. Passalacqua
G. Passaleva
D. Passaro
A. Pastore
M. Patel
J. Patoc
C. Patrignani
C.J. Pawley
A. Pellegrino
M. Pepe Altarelli
S. Perazzini
D. Pereima
A. Pereiro Castro
P. Perret
A. Perro
K. Petridis
A. Petrolini
S. Petrucci
J. P. Pfaller
H. Pham
L. Pica
M. Piccini
B. Pietrzyk
G. Pietrzyk
D. Pinci
F. Pisani
M. Pizzichemi
V. Placinta
M. Plo Casasus
F. Polci
M. Poli Lener
A. Poluektov
N. Polukhina
I. Polyakov
E. Polycarpo
S. Ponce
D. Popov
S. Poslavskii
K. Prasanth
C. Prouve
V. Pugatch
G. Punzi
W. Qian
N. Qin
S. Qu
R. Quagliani
R.I. Rabadan Trejo
J.H. Rademacker
M. Rama
M. Ramírez García
M. Ramos Pernas
M.S. Rangel
F. Ratnikov
G. Raven
M. Rebollo De Miguel
F. Redi
J. Reich
F. Reiss
Z. Ren
P.K. Resmi
R. Ribatti
G. R. Ricart
D. Riccardi
S. Ricciardi
K. Richardson
M. Richardson-Slipper
K. Rinnert
P. Robbe
G. Robertson
E. Rodrigues
E. Rodriguez Fernandez
J.A. Rodriguez Lopez
E. Rodriguez Rodriguez
A. Rogovskiy
D.L. Rolf
A. Rollings
P. Roloff
V. Romanovskiy
M. Romero Lamas
A. Romero Vidal
G. Romolini
F. Ronchetti
M. Rotondo
S. R. Roy
M.S. Rudolph
T. Ruf
M. Ruiz Diaz
R.A. Ruiz Fernandez
J. Ruiz Vidal
A. Ryzhikov
J. Ryzka
J.J. Saborido Silva
R. Sadek
N. Sagidova
N. Sahoo
B. Saitta
M. Salomoni
C. Sanchez Gras
I. Sanderswood
R. Santacesaria
C. Santamarina Rios
M. Santimaria
L. Santoro
E. Santovetti
A. Saputi
D. Saranin
G. Sarpis
M. Sarpis
A. Sarti
C. Satriano
A. Satta
M. Saur
D. Savrina
H. Sazak
L.G. Scantlebury Smead
A. Scarabotto
S. Schael
S. Scherl
A. M. Schertz
M. Schiller
H. Schindler
M. Schmelling
B. Schmidt
S. Schmitt
H. Schmitz,
O. Schneider
A. Schopper
N. Schulte
S. Schulte
M.H. Schune
R. Schwemmer
G. Schwering
B. Sciascia
A. Sciuccati
S. Sellam
A. Semennikov
M. Senghi Soares
A. Sergi
N. Serra
L. Sestini
A. Seuthe
Y. Shang
D.M. Shangase
M. Shapkin
R. S. Sharma
I. Shchemerov
L. Shchutska
T. Shears
L. Shekhtman
Z. Shen
S. Sheng
V. Shevchenko
B. Shi
E.B. Shields
Y. Shimizu
E. Shmanin
R. Shorkin
J.D. Shupperd
R. Silva Coutinho
G. Simi
S. Simone
N. Skidmore
R. Skuza
T. Skwarnicki
M.W. Slater
J.C. Smallwood
E. Smith
K. Smith
M. Smith
A. Snoch
L. Soares Lavra
M.D. Sokoloff
F.J.P. Soler
A. Solomin
A. Solovev
I. Solovyev
R. Song
Y. Song
Y. Song
Y. S. Song
F.L. Souza De Almeida
B. Souza De Paula
E. Spadaro Norella
E. Spedicato
J.G. Speer
E. Spiridenkov,
P. Spradlin
V. Sriskaran
F. Stagni
M. Stahl
S. Stahl
S. Stanislaus
E.N. Stein
O. Steinkamp
O. Stenyakin,
H. Stevens
D. Strekalina
Y. Su
F. Suljik
J. Sun
L. Sun
Y. Sun
P.N. Swallow
F. Swystun
A. Szabelski
T. Szumlak
M. Szymanski
Y. Tan
S. Taneja
M.D. Tat
A. Terentev
F. Terzuoli
F. Teubert
E. Thomas
D.J.D. Thompson
H. Tilquin
V. Tisserand
S. T’Jampens
M. Tobin
L. Tomassetti
G. Tonani
X. Tong
D. Torres Machado
L. Toscano
D.Y. Tou
C. Trippl
G. Tuci
N. Tuning
L.H. Uecker
A. Ukleja
D.J. Unverzagt
E. Ursov
A. Usachov
A. Ustyuzhanin
U. Uwer
V. Vagnoni
A. Valassi
G. Valenti
N. Valls Canudas
H. Van Hecke
E. van Herwijnen
C.B. Van Hulse
R. Van Laak
M. van Veghel
R. Vazquez Gomez
P. Vazquez Regueiro
C. Vázquez Sierra
S. Vecchi
J.J. Velthuis
M. Veltri
A. Venkateswaran
M. Vesterinen
M. Vieites Diaz
X. Vilasis-Cardona
E. Vilella Figueras
A. Villa
P. Vincent
F.C. Volle
D. vom Bruch
V. Vorobyev,
N. Voropaev
K. Vos
G. Vouters,
C. Vrahas
J. Walsh
E.J. Walton
G. Wan
C. Wang
G. Wang
J. Wang
J. Wang
J. Wang
J. Wang
M. Wang
N. W. Wang
R. Wang
X. Wang
X. W. Wang
Y. Wang
Z. Wang
Z. Wang
Z. Wang
J.A. Ward
M. Waterlaat,
N.K. Watson
D. Websdale
Y. Wei
B.D.C. Westhenry
C. White
D.J. White
M. Whitehead
A.R. Wiederhold
D. Wiedner
G. Wilkinson
M.K. Wilkinson
M. Williams
M.R.J. Williams
R. Williams
F.F. Wilson
W. Wislicki
M. Witek
L. Witola
C.P. Wong
G. Wormser
S.A. Wotton
H. Wu
J. Wu
Y. Wu
K. Wyllie
S. Xian,
Z. Xiang
Y. Xie
A. Xu
J. Xu
L. Xu
L. Xu
M. Xu
Z. Xu
Z. Xu
Z. Xu
D. Yang
S. Yang
X. Yang
Y. Yang
Z. Yang
Z. Yang
V. Yeroshenko
H. Yeung
H. Yin
C. Y. Yu
J. Yu
X. Yuan
E. Zaffaroni
M. Zavertyaev
M. Zdybal
M. Zeng
C. Zhang
D. Zhang
J. Zhang
L. Zhang
S. Zhang
S. Zhang
Y. Zhang
Y. Z. Zhang
Y. Zhao
A. Zharkova
A. Zhelezov
X. Z. Zheng
Y. Zheng
T. Zhou
X. Zhou
Y. Zhou
V. Zhovkovska
L. Z. Zhu
X. Zhu
X. Zhu
V. Zhukov
J. Zhuo
Q. Zou
D. Zuliani
G. Zunica
School of Physics and Astronomy, Monash University, Melbourne, Australia
Centro Brasileiro de Pesquisas Físicas (CBPF), Rio de Janeiro, Brazil
Universidade Federal do Rio de Janeiro (UFRJ), Rio de Janeiro, Brazil
Center for High Energy Physics, Tsinghua University, Beijing, China
Institute Of High Energy Physics (IHEP), Beijing, China
School of Physics State Key Laboratory of Nuclear Physics and Technology, Peking University, Beijing, China
University of Chinese Academy of Sciences, Beijing, China
Institute of Particle Physics, Central China Normal University, Wuhan, Hubei, China
Consejo Nacional de Rectores (CONARE), San Jose, Costa Rica
Université Savoie Mont Blanc, CNRS, IN2P3-LAPP, Annecy, France
Université Clermont Auvergne, CNRS/IN2P3, LPC, Clermont-Ferrand, France
Aix Marseille Univ, CNRS/IN2P3, CPPM, Marseille, France
Université Paris-Saclay, CNRS/IN2P3, IJCLab, Orsay, France
Laboratoire Leprince-Ringuet, CNRS/IN2P3, Ecole Polytechnique, Institut Polytechnique de Paris, Palaiseau, France
LPNHE, Sorbonne Université, Paris Diderot Sorbonne Paris Cité, CNRS/IN2P3, Paris, France
I. Physikalisches Institut, RWTH Aachen University, Aachen, Germany
Fakultät Physik, Technische Universität Dortmund, Dortmund, Germany
Max-Planck-Institut für Kernphysik (MPIK), Heidelberg, Germany
Physikalisches Institut, Ruprecht-Karls-Universität Heidelberg, Heidelberg, Germany
School of Physics, University College Dublin, Dublin, Ireland
INFN Sezione di Bari, Bari, Italy
INFN Sezione di Bologna, Bologna, Italy
INFN Sezione di Ferrara, Ferrara, Italy
INFN Sezione di Firenze, Firenze, Italy
INFN Laboratori Nazionali di Frascati, Frascati, Italy
INFN Sezione di Genova, Genova, Italy
INFN Sezione di Milano, Milano, Italy
INFN Sezione di Milano-Bicocca, Milano, Italy
INFN Sezione di Cagliari, Monserrato, Italy
Università degli Studi di Padova, Università e INFN, Padova, Padova, Italy
INFN Sezione di Perugia, Perugia, Italy
INFN Sezione di Pisa, Pisa, Italy
INFN Sezione di Roma La Sapienza, Roma, Italy
INFN Sezione di Roma Tor Vergata, Roma, Italy
Nikhef National Institute for Subatomic Physics, Amsterdam, Netherlands
Nikhef National Institute for Subatomic Physics and VU University Amsterdam, Amsterdam, Netherlands
AGH - University of Science and Technology, Faculty of Physics and Applied Computer Science, Kraków, Poland
Henryk Niewodniczanski Institute of Nuclear Physics Polish Academy of Sciences, Kraków, Poland
National Center for Nuclear Research (NCBJ), Warsaw, Poland
Horia Hulubei National Institute of Physics and Nuclear Engineering, Bucharest-Magurele, Romania
Affiliated with an institute covered by a cooperation agreement with CERN
DS4DS, La Salle, Universitat Ramon Llull, Barcelona, Spain
ICCUB, Universitat de Barcelona, Barcelona, Spain
Instituto Galego de Física de Altas Enerxías (IGFAE), Universidade de Santiago de Compostela, Santiago de Compostela, Spain
Instituto de Fisica Corpuscular, Centro Mixto Universidad de Valencia - CSIC, Valencia, Spain
European Organization for Nuclear Research (CERN), Geneva, Switzerland
Institute of Physics, Ecole Polytechnique Fédérale de Lausanne (EPFL), Lausanne, Switzerland
Physik-Institut, Universität Zürich, Zürich, Switzerland
NSC Kharkiv Institute of Physics and Technology (NSC KIPT), Kharkiv, Ukraine
Institute for Nuclear Research of the National Academy of Sciences (KINR), Kyiv, Ukraine
University of Birmingham, Birmingham, United Kingdom
H.H. Wills Physics Laboratory, University of Bristol, Bristol, United Kingdom
Cavendish Laboratory, University of Cambridge, Cambridge, United Kingdom
Department of Physics, University of Warwick, Coventry, United Kingdom
STFC Rutherford Appleton Laboratory, Didcot, United Kingdom
School of Physics and Astronomy, University of Edinburgh, Edinburgh, United Kingdom
School of Physics and Astronomy, University of Glasgow, Glasgow, United Kingdom
Oliver Lodge Laboratory, University of Liverpool, Liverpool, United Kingdom
Imperial College London, London, United Kingdom
Department of Physics and Astronomy, University of Manchester, Manchester, United Kingdom
Department of Physics, University of Oxford, Oxford, United Kingdom
Massachusetts Institute of Technology, Cambridge, MA, United States
University of Cincinnati, Cincinnati, OH, United States
University of Maryland, College Park, MD, United States
Los Alamos National Laboratory (LANL), Los Alamos, NM, United States
Syracuse University, Syracuse, NY, United States
Pontifícia Universidade Católica do Rio de Janeiro (PUC-Rio), Rio de Janeiro, Brazil, associated to
School of Physics and Electronics, Hunan University, Changsha City, China, associated to
Guangdong Provincial Key Laboratory of Nuclear Science, Guangdong-Hong Kong Joint Laboratory of Quantum Matter, Institute of Quantum Matter, South China Normal University, Guangzhou, China, associated to
Lanzhou University, Lanzhou, China, associated to
School of Physics and Technology, Wuhan University, Wuhan, China, associated to
[…]

### v2

@@ I. Kostiuk35
A. Kozachuk41
P. Kravchenko41
L. Kravchuk41
M. Kreps54
S. Kretzschmar16
P. Krokovny41
W. Krupa66
W. Krzemien39
J. Kubat19,
S. Kubis77
W. Kucewicz38
M. Kucharczyk38
V. Kudryavtsev41
E. Kulikova41
A. Kupsc78
B. K. Kutsenko12
D. Lacarrere46
A. Lai29
A. Lampis29
D. Lancierini48
C. Landesa Gomez44
J.J. Lane1
R. Lane52
C. Langenbruch19
J. Langer17
O. Lantwin41
T. Latham54
F. Lazzari32,s
C. Lazzeroni51
R. Le Gac12
S.H. Lee79
R. Lefèvre11
A. Leflat41
S. Legotin41
M. Lehuraux54
O. Leroy12
T. Lesiak38
B. Leverington19
A. Li4
H. Li69
K. Li8
L. Li60
P. Li46
P.-R. Li70
S. Li8
T. Li5,d
T. Li69
Y. Li8,
Y. Li5
Z. Li66
Z. Lian4
X. Liang66
C. Lin7
T. Lin55
R. Lindner46
V. Lisovskyi47
R. Litvinov29,j
F. L. Liu1
G. Liu69
K. Liu70
Q. Liu7
S. Liu5,7
Y. Liu56
Y. Liu70,
Y. L. Liu59
A. Lobo Salvia43
A. Loi29
J. Lomba Castro44
T. Long53
J.H. Lopes3
A. Lopez Huertas43
S. López Soliño44
G.H. Lovell53
C. Lucarelli24,l
D. Lucchesi30,p
S. Luchuk41
M. Lucio Martinez76
V. Lukashenko35,50
Y. Luo6
A. Lupato30
E. Luppi23,k
K. Lynch20
X.-R. Lyu7
G. M. Ma4
R. Ma7
S. Maccolini17
F. Machefert13
F. Maciuc40
B. M. Mack66
I. Mackay61
L. M. Mackey66
L.R. Madhan Mohan53
M. M. Madurai51
A. Maevskiy41
D. Magdalinski35
V. Mahajan54
D. Maisuzenko41
M.W. Majewski37,
J.J. Malczewski38
S. Malde61
B. Malecki38,46
L. Malentacca46,
A. Malinin41
T. Maltsev41
G. Manca29,j
G. Mancinelli12
C. Mancuso27,13,n
R. Manera Escalero43,
D. Manuzzi22
D. Marangotto27,n
J.F. Marchand10
R. Marchevski47
U. Marconi22
S. Mariani46
C. Marin Benito43
J. Marks19
A.M. Marshall52
P.J. Marshall58,
G. Martelli31,q
G. Martellotti33
L. Martinazzoli46
M. Martinelli28,o
D. Martinez Santos44
F. Martinez Vidal45
A. Massafferri2
M. Materok16
R. Matev46
A. Mathad48
V. Matiunin41
C. Matteuzzi66
K.R. Mattioli14
A. Mauri59
E. Maurice14
J. Mauricio43
P. Mayencourt47
M. Mazurek46
M. McCann59
L. Mcconnell20
T.H. McGrath60
N.T. McHugh57
A. McNab60
R. McNulty20
B. Meadows63
G. Meier17
D. Melnychuk39
M. Merk35,76
A. Merli27,n
L. Meyer Garcia3
D. Miao5,7
H. Miao7
M. Mikhasenko73,f
D.A. Milanes72
A. Minotti28,o
E. Minucci66
T. Miralles11
S.E. Mitchell56
B. Mitreska17
D.S. Mitzel17
A. Modak55
A. Mödden 17
R.A. Mohammed61
R.D. Moise16
S. Mokhnenko41
T. Mombächer46
M. Monk54,1
I.A. Monroy72
S. Monteil11
A. Morcillo Gomez44
G. Morello25
M.J. Morello32,r
M.P. Morgenthaler19
A.B. Morris46
A.G. Morris12
R. Mountain66
H. Mu4
Z. M. Mu6
E. Muhammad54
F. Muheim56
M. Mulder75
K. Müller48
F. Muñoz-Rojas9
R. Murta59
P. Naik58
T. Nakada47
R. Nandakumar55
T. Nanut46
I. Nasteva3
M. Needham56
N. Neri27,n
S. Neubert73
N. Neufeld46
P. Neustroev41,
R. Newcombe59,
J. Nicolini17,13
D. Nicotra76
E.M. Niel47
N. Nikitin41
P. Nogga73,
N.S. Nolte62
C. Normand10,29
J. Novoa Fernandez44
G. Nowak63
C. Nunez79
H. N. Nur57
A. Oblakowska-Mucha37
V. Obraztsov41
T. Oeser16
S. Okamura23,k,46
R. Oldeman29,j
F. Oliva56
M. Olocco17
C.J.G. Onderwater76
R.H. O’Neil56
J.M. Otalora Goicochea3
T. Ovsiannikova41
P. Owen48
A. Oyanguren45
O. Ozcelik56
K.O. Padeken73
B. Pagare54
P.R. Pais19
T. Pajero61
A. Palano21
M. Palutan25
G. Panshin41
L. Paolucci54
A. Papanestis55
M. Pappagallo21,h
L.L. Pappalardo23,k
C. Pappenheimer63
C. Parkes60
B. Passalacqua23,k
G. Passaleva24
D. Passaro32,r
A. Pastore21
M. Patel59
J. Patoc61
C. Patrignani22,i
C.J. Pawley76
A. Pellegrino35
M. Pepe Altarelli25
S. Perazzini22
D. Pereima41
A. Pereiro Castro44
P. Perret11
A. Perro46
K. Petridis52
A. Petrolini26,m
S. Petrucci56
J. P. Pfaller63
H. Pham66
L. Pica32,r
M. Piccini31
B. Pietrzyk10
G. Pietrzyk13
D. Pinci33
F. Pisani46
M. Pizzichemi28,o
V. Placinta40
M. Plo Casasus44
F. Polci15,46
M. Poli Lener25
A. Poluektov12
N. Polukhina41
I. Polyakov46
E. Polycarpo3
S. Ponce46
D. Popov7
S. Poslavskii41
K. Prasanth38
C. Prouve44
V. Pugatch50
G. Punzi32,s
W. Qian7
N. Qin4
S. Qu4
R. Quagliani47
R.I. Rabadan Trejo54
J.H. Rademacker52
M. Rama32
M. Ramírez García79
M. Ramos Pernas54
M.S. Rangel3
F. Ratnikov41
G. Raven36
M. Rebollo De Miguel45
F. Redi46
J. Reich52
F. Reiss60
Z. Ren7
P.K. Resmi61
R. Ribatti32,r
G. R. Ricart14,80
D. Riccardi32,r
S. Ricciardi55
K. Richardson62
M. Richardson-Slipper56
K. Rinnert58
P. Robbe13
G. Robertson57
E. Rodrigues58,46
E. Rodriguez Fernandez44
J.A. Rodriguez Lopez72
E. Rodriguez Rodriguez44
A. Rogovskiy55
D.L. Rolf46
A. Rollings61
P. Roloff46
V. Romanovskiy41
M. Romero Lamas44
A. Romero Vidal44
G. Romolini23
F. Ronchetti47
M. Rotondo25
S. R. Roy19
M.S. Rudolph66
T. Ruf46
M. Ruiz Diaz19
R.A. Ruiz Fernandez44
J. Ruiz Vidal78,z
A. Ryzhikov41
J. Ryzka37
J.J. Saborido Silva44
R. Sadek14
N. Sagidova41
N. Sahoo51
B. Saitta29,j
M. Salomoni28,o
C. Sanchez Gras35
I. Sanderswood45
R. Santacesaria33
C. Santamarina Rios44
M. Santimaria25
L. Santoro 2
E. Santovetti34
A. Saputi23,46
D. Saranin41
G. Sarpis56
M. Sarpis73
A. Sarti33
C. Satriano33,t
A. Satta34
M. Saur6
D. Savrina41
H. Sazak11
L.G. Scantlebury Smead61
A. Scarabotto15
S. Schael16
S. Scherl58
A. M. Schertz74
M. Schiller57
H. Schindler46
M. Schmelling18
B. Schmidt46
S. Schmitt16
H. Schmitz73,
O. Schneider47
A. Schopper46
N. Schulte17
S. Schulte47
M.H. Schune13
R. Schwemmer46
G. Schwering16
B. Sciascia25
A. Sciuccati46
S. Sellam44
A. Semennikov41
M. Senghi Soares36
A. Sergi26,m
N. Serra48,46
L. Sestini30
A. Seuthe17
Y. Shang6
D.M. Shangase79
M. Shapkin41
R. S. Sharma66
I. Shchemerov41
L. Shchutska47
T. Shears58
L. Shekhtman41
Z. Shen6
S. Sheng5,7
V. Shevchenko41
B. Shi7
E.B. Shields28,o
Y. Shimizu13
E. Shmanin41
R. Shorkin41
J.D. Shupperd66
R. Silva Coutinho66
G. Simi30
S. Simone21,h
N. Skidmore54
R. Skuza19
T. Skwarnicki66
M.W. Slater51
J.C. Smallwood61
E. Smith62
K. Smith65
M. Smith59
A. Snoch35
L. Soares Lavra56
M.D. Sokoloff63
F.J.P. Soler57
A. Solomin41,52
A. Solovev41
I. Solovyev41
R. Song1
Y. Song47
Y. Song4
Y. S. Song6
F.L. Souza De Almeida66
B. Souza De Paula3
E. Spadaro Norella27,n
E. Spedicato22
J.G. Speer17
E. Spiridenkov41,
P. Spradlin57
V. Sriskaran46
F. Stagni46
M. Stahl46
S. Stahl46
S. Stanislaus61
E.N. Stein46
O. Steinkamp48
O. Stenyakin41,
H. Stevens17
D. Strekalina41
Y. Su7
F. Suljik61
J. Sun29
L. Sun71
Y. Sun64
P.N. Swallow51
F. Swystun54
A. Szabelski39
T. Szumlak37
M. Szymanski46
Y. Tan4
S. Taneja60
M.D. Tat61
A. Terentev48
F. Terzuoli32,v
F. Teubert46
E. Thomas46
D.J.D. Thompson51
H. Tilquin59
V. Tisserand11
S. T’Jampens10
M. Tobin5
L. Tomassetti23,k
G. Tonani27,n,46
X. Tong6
D. Torres Machado2
L. Toscano17
D.Y. Tou4
C. Trippl42
G. Tuci19
N. Tuning35
L.H. Uecker19
A. Ukleja37
D.J. Unverzagt19
E. Ursov41
A. Usachov36
A. Ustyuzhanin41
U. Uwer19
V. Vagnoni22
A. Valassi46
G. Valenti22
N. Valls Canudas42
H. Van Hecke65
E. van Herwijnen59
C.B. Van Hulse44,x
R. Van Laak47
M. van Veghel35
R. Vazquez Gomez43
P. Vazquez Regueiro44
C. Vázquez Sierra44
S. Vecchi23
J.J. Velthuis52
M. Veltri24,w
A. Venkateswaran47
M. Vesterinen54
M. Vieites Diaz46
X. Vilasis-Cardona42
E. Vilella Figueras58
A. Villa22
P. Vincent15
F.C. Volle13
D. vom Bruch12
V. Vorobyev41,
N. Voropaev41
K. Vos76
G. Vouters10,
C. Vrahas56
J. Walsh32
E.J. Walton1
G. Wan6
C. Wang19
G. Wang8
J. Wang6
J. Wang5
J. Wang4
J. Wang71
M. Wang27
N. W. Wang7
R. Wang52
X. Wang69
X. W. Wang59
Y. Wang8
Z. Wang13
Z. Wang4
Z. Wang7
J.A. Ward54,1
M. Waterlaat46,
N.K. Watson51
D. Websdale59
Y. Wei6
B.D.C. Westhenry52
C. White54
D.J. White60
M. Whitehead57
A.R. Wiederhold54
D. Wiedner17
G. Wilkinson61
M.K. Wilkinson63
M. Williams62
M.R.J. Williams56
R. Williams53
F.F. Wilson55
W. Wislicki39
M. Witek38
L. Witola19
C.P. Wong65
G. Wormser13
S.A. Wotton53
H. Wu66
J. Wu8
Y. Wu6
K. Wyllie46
S. Xian69,
Z. Xiang5
Y. Xie8
A. Xu32
J. Xu7
L. Xu4
L. Xu4
M. Xu54
Z. Xu11
Z. Xu7
Z. Xu5
D. Yang4
S. Yang7
X. Yang6
Y. Yang26,m
Z. Yang6
Z. Yang64
V. Yeroshenko13
H. Yeung60
H. Yin8
C. Y. Yu6
J. Yu68
X. Yuan5
E. Zaffaroni47
M. Zavertyaev18
M. Zdybal38
M. Zeng4
C. Zhang6
D. Zhang8
J. Zhang7
L. Zhang4
S. Zhang68
S. Zhang6
Y. Zhang6
Y. Z. Zhang4
Y. Zhao19
A. Zharkova41
A. Zhelezov19
X. Z. Zheng4
Y. Zheng7
T. Zhou6
X. Zhou8
Y. Zhou7
V. Zhovkovska54
L. Z. Zhu7
X. Zhu4
X. Zhu8
V. Zhukov16,41
J. Zhuo45
Q. Zou5,7
D. Zuliani30
G. Zunica60
1School of Physics and Astronomy, Monash University, Melbourne, Australia
2Centro Brasileiro de Pesquisas Físicas (CBPF), Rio de Janeiro, Brazil
3Universidade Federal do Rio de Janeiro (UFRJ), Rio de Janeiro, Brazil
4Center for High Energy Physics, Tsinghua University, Beijing, China
5Institute Of High Energy Physics (IHEP), Beijing, China
6School of Physics State Key Laboratory of Nuclear Physics and Technology, Peking University, Beijing, China
7University of Chinese Academy of Sciences, Beijing, China
8Institute of Particle Physics, Central China Normal University, Wuhan, Hubei, China
9Consejo Nacional de Rectores (CONARE), San Jose, Costa Rica
10Université Savoie Mont Blanc, CNRS, IN2P3-LAPP, Annecy, France
11Université Clermont Auvergne, CNRS/IN2P3, LPC, Clermont-Ferrand, France
12Aix Marseille Univ, CNRS/IN2P3, CPPM, Marseille, France
13Université Paris-Saclay, CNRS/IN2P3, IJCLab, Orsay, France
14Laboratoire Leprince-Ringuet, CNRS/IN2P3, Ecole Polytechnique, Institut Polytechnique de Paris, Palaiseau, France
15LPNHE, Sorbonne Université, Paris Diderot Sorbonne Paris Cité, CNRS/IN2P3, Paris, France
16I. Physikalisches Institut, RWTH Aachen University, Aachen, Germany
17Fakultät Physik, Technische Universität Dortmund, Dortmund, Germany
18Max-Planck-Institut für Kernphysik (MPIK), Heidelberg, Germany
19Physikalisches Institut, Ruprecht-Karls-Universität Heidelberg, Heidelberg, Germany
20School of Physics, University College Dublin, Dublin, Ireland
21INFN Sezione di Bari, Bari, Italy
22INFN Sezione di Bologna, Bologna, Italy
23INFN Sezione di Ferrara, Ferrara, Italy
24INFN Sezione di Firenze, Firenze, Italy
25INFN Laboratori Nazionali di Frascati, Frascati, Italy
26INFN Sezione di Genova, Genova, Italy
27INFN Sezione di Milano, Milano, Italy
28INFN Sezione di Milano-Bicocca, Milano, Italy
29INFN Sezione di Cagliari, Monserrato, Italy
30Università degli Studi di Padova, Università e INFN, Padova, Padova, Italy
31INFN Sezione di Perugia, Perugia, Italy
32INFN Sezione di Pisa, Pisa, Italy
33INFN Sezione di Roma La Sapienza, Roma, Italy
34INFN Sezione di Roma Tor Vergata, Roma, Italy
35Nikhef National Institute for Subatomic Physics, Amsterdam, Netherlands
36Nikhef National Institute for Subatomic Physics and VU University Amsterdam, Amsterdam, Netherlands
37AGH - University of Science and Technology, Faculty of Physics and Applied Computer Science, Kraków, Poland
38Henryk Niewodniczanski Institute of Nuclear Physics Polish Academy of Sciences, Kraków, Poland
39National Center for Nuclear Research (NCBJ), Warsaw, Poland
40Horia Hulubei National Institute of Physics and Nuclear Engineering, Bucharest-Magurele, Romania
41Affiliated with an institute covered by a cooperation agreement with CERN
42DS4DS, La Salle, Universitat Ramon Llull, Barcelona, Spain
43ICCUB, Universitat de Barcelona, Barcelona, Spain
44Instituto Galego de Física de Altas Enerxías (IGFAE), Universidade de Santiago de Compostela, Santiago de Compostela, Spain
45Instituto de Fisica Corpuscular, Centro Mixto Universidad de Valencia - CSIC, Valencia, Spain
46European Organization for Nuclear Research (CERN), Geneva, Switzerland
47Institute of Physics, Ecole Polytechnique Fédérale de Lausanne (EPFL), Lausanne, Switzerland
48Physik-Institut, Universität Zürich, Zürich, Switzerland
49NSC Kharkiv Institute of Physics and Technology (NSC KIPT), Kharkiv, Ukraine
50Institute for Nuclear Research of the National Academy of Sciences (KINR), Kyiv, Ukraine
51University of Birmingham, Birmingham, United Kingdom
52H.H. Wills Physics Laboratory, University of Bristol, Bristol, United Kingdom
53Cavendish Laboratory, University of Cambridge, Cambridge, United Kingdom
54Department of Physics, University of Warwick, Coventry, United Kingdom
55STFC Rutherford Appleton Laboratory, Didcot, United Kingdom
56School of Physics and Astronomy, University of Edinburgh, Edinburgh, United Kingdom
57School of Physics and Astronomy, University of Glasgow, Glasgow, United Kingdom
58Oliver Lodge Laboratory, University of Liverpool, Liverpool, United Kingdom
59Imperial College London, London, United Kingdom
60Department of Physics and Astronomy, University of Manchester, Manchester, United Kingdom
61Department of Physics, University of Oxford, Oxford, United Kingdom
62Massachusetts Institute of Technology, Cambridge, MA, United States
63University of Cincinnati, Cincinnati, OH, United States
64University of Maryland, College Park, MD, United States
65Los Alamos National Laboratory (LANL), Los Alamos, NM, United States
66Syracuse University, Syracuse, NY, United States
67Pontifícia Universidade Católica do Rio de Janeiro (PUC-Rio), Rio de Janeiro, Brazil, associated to 3
68School of Physics and Electronics, Hunan University, Changsha City, China, associated to 8
69Guangdong Provincial Key Laboratory of Nuclear Science, Guangdong-Hong Kong Joint Laboratory of Quantum Matter, Institute of Quantum Matter, South China Normal University, Guangzhou, China, associated to 4
70Lanzhou University, Lanzhou, China, associated to 5
[…]